# Understanding Bias Reinforcement in LLM Agents Debate

### v4

@@ B.1 Game Prompts
#Task:
Based on the current state of the game, decide which pin(s) you will take on this turn.
Table 13: Chomp game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a game of Chomp.
#Objective:
Your goal is to force your opponent to take the top-left corner of the grid (position (0, 0)).
#Game Rule:
1. The game is played on a square grid.
2. On your turn, you select a position (row, col).
4. The player forced to select (0, 0) loses.
#Current State:
The grid is represented as a binary matrix, where ’1’ means the position is still available, and ’0’ means it is removed: {remaining_grid}
#Task:
Based on the current state of the grid, decide which position (row, col) you will select.
Table 14: Corner-Queen game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a Corner-Queen game.
#Objective:
Move the queen so that you are the first to place it on the lower-left corner square.
#Game Rule:
@@
2. Coordinates use zero-based indices [row, col]. Row 0 is the top row; Col 0 is the leftmost column. Valid ranges: row $\in[0,\ \texttt{board\_height}-1]$, col $\in[0,\ \texttt{board\_width}-1]$.
3. From the current position $[r,c]$ the queen may move to **one** of:  (a) left: $r,\ c^{\prime}]$ with $c^{\prime}<c$;
 (b) down: $r^{\prime},\ c]$ with $r^{\prime}>r$;
 (c) left-down diagonal: [r+d,cd][r+d,\ c-d] with $d>0$.
@@
#Current State:
Current position: [row = {r}, col = {c}].
#Task:
Based on the current state, decide the next move [row,col][row,\ col].
B.2 Prompts for basic reasoning
B.2.1 Standard, ReAct & CoT Prompts
In our evaluation of LLMs within the MetaNIM Arena, we compare two key prompting techniques: Standard Prompting and Chain-of-Thought (CoT) Prompting. The distinction between these approaches significantly impacts the model’s reasoning and decision-making process.
Standard Prompting (Table 15)
Standard prompting provides a direct task description, outlining the game rules, current state, and the required decision. The model is expected to generate only an action to determine the best move. This method is cheap and efficient but often leads to suboptimal decisions, as the model may fail to make proper reasoning before selecting the action.
ReAct Prompting (Table 16)
ReAct prompting provides a direct task description, outlining the game rules, current state, and the required decision. The model is expected to generate an action with proper explicit reasoning steps to determine the best move. This method is efficient but often leads to suboptimal decisions, as the model may fail to retrieve and apply deeper strategic reasoning.
Chain-of-Thought (CoT) Prompting (Table 17)
CoT prompting extends the standard prompt by explicitly instructing the model to think step-by-step before making a decision. By guiding the model through an explicit reasoning process, CoT enables it to break down the problem, consider strategic implications, and refine its choices before committing to an action. This often leads to improved decision-making, particularly in multi-step strategic environments where deeper reasoning is required.
Key Difference and Impact
As illustrated in Table 17, the only modification in the CoT prompt is the addition of a simple directive: “Let’s think step-by-step. What is the best move for you?” This small change significantly alters the model’s reasoning trajectory, encouraging more structured and strategic decision-making. Our experimental results (detailed in §3) confirm that CoT prompting leads to a measurable improvement in decision accuracy, particularly in complex scenarios where retrieving and applying prior knowledge is essential.
By leveraging CoT, we can enhance the model’s ability to explain its decisions, mitigate biases, and adapt more effectively to adversarial settings. However, as we further discuss in Experiment Section, CoT has limitations to leverage the strategic reasoining well in our proposed environment, necessitating additional mechanisms to further enhance strategic reasoning.
Table 15: Standard Prompt in NIM
#Game Role:
You are {agent[’name’]}, a participant in a game of Nim variants.
#Objective:
Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins.
#Game Rule:
There is a single pile of items. You can take between 1 and {max_take} items on your turn.
#Current State:
There are {remaining_items} items remaining in the pile.
#Task:
Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn.
Output Format:
The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and:
action: integer // This is an action you take. Only integer between 1 and 3.
Table 16: ReAct Prompt in NIM
#Game Role:
You are {agent[’name’]}, a participant in a game of Nim variants.
#Objective:
Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins.
#Game Rule:
There is a single pile of items. You can take between 1 and {max_take} items on your turn.
#Current State:
There are {remaining_items} items remaining in the pile.
#Task:
Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn.
Output Format:
The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and:
reasoning: string // This is the reason for the action
action: integer // This is an action you take based on the reasoning. Only integer between 1 and 3.
Table 17: CoT Prompt in NIM
#Game Role:
You are {agent[’name’]}, a participant in a game of Nim variants.
#Objective:
Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins.
#Game Rule:
There is a single pile of items. You can take between 1 and {max_take} items on your turn.
#Current State:
There are {remaining_items} items remaining in the pile.
#Task:
Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn.
Output Format:
The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and:
reasoning: string // This is the reason for the action
action: integer // This is an action you take based on the reasoning. Only integer between 1 and 3.
Let’s think step-by-step. What is the best move for you?
Prompting Strategy for Opponent Modeling
@@ B.2.1 Standard, ReAct & CoT Prompts
Appendix C DReaMAD (-): Structured Prompt Optimization without Debate
While the full DReaMAD framework integrates multi-agent debate to refine strategic reasoning, its core prompting methodology—excluding debate—remains a powerful mechanism for enhancing decision-making. This streamlined version, DReaMAD (-), focuses on three key stages to systematically extract and refine strategic knowledge, improving reasoning diversity and mitigating bias. We present the DReaMAD prompt as in Table 18.
1. Game Situation Reinterpretation
The first step involves extracting fundamental game principles from the standard prompt. The model is tasked with identifying key elements, such as:
  • Game Definition: The nature of the game and its mechanics.
  • Winning Condition: The criteria for victory.
  • Move Constraints: The permissible actions per turn.
This step ensures that the model builds a structured understanding of the strategic environment before making decisions.
2. General Strategy Formulation
After extracting the core game elements, the model derives a generalized winning strategy applicable to various game states. It generates:
  • State Evaluation: How to assess the game state at any given turn.
  • Winning Strategy: The optimal decision-making framework for victory.
  • Endgame Tactics: Best strategies in near-win scenarios.
This formulation helps structure the model’s reasoning beyond the immediate game context, fostering more strategic foresight.
3. Perspective Diversification
Finally, the model refines the original prompt using the extracted strategic knowledge. This process introduces structured variations to the initial prompt to encourage diverse reasoning, rather than reinforcing a singular bias. The self-refined prompt:
  • Guides decision-making explicitly.
  • Prioritizes winning strategies.
  • Encourages logical, step-by-step reasoning.
This structured refinement ensures that LLMs adopt distinct strategic viewpoints even without external debate, improving their adaptability and robustness in adversarial environments.
By systematically structuring knowledge retrieval and refining prompts, DReaMAD (-) enhances strategic reasoning as illustrated in Figure 10, this approach strengthens the model’s ability to retrieve and apply prior knowledge effectively, offering a scalable solution for improving LLM-based decision-making.
Table 18: DReaMAD prompt design before debate
#Game Situation Reinterpretation:
game_prompt :
Below is a game description. Extract the key information.
Game Description: {current state}
### Format response as:
game definition: string // What is the definition of this game?
winning condition: string // How to win the game.
move constraints: string // What actions are allowed per turn.
#General Strategy Formulation:
strategy_prompt :
Based on the game information below, derive the general winning strategy in this game
Game: {game definition}
Winning Condition: {winning condition}
Move Constraints: {move constraints}
Current State: {current state in very short}
### Format response as:
state evaluation: string // How to assess the game state.
winning strategy: string // Winning strategy in this turn to win this game.
endgame tactics: string // Best strategy in a near-win situation.
#Perspective Diversification:
Refine the initial game prompt to improve decision-making based on the Game and Strategy Information. Initial Prompt: {given initial prompt}
Game and Strategy Information:
Game: {game definition}
Strategy:
- State Evaluation: {state evaluation}
- Winning Strategy: {winning strategy}
- Endgame Tactics: {endgame tactics}
### Format response as:
optimized prompt: string // The refined prompt that clearly directs decision-making.
Appendix D A Word Change in Prompts Lead to Different Output
Figure 8: According the word Fibonacci usage, the reasoning and the performance differs.
D.1 Prompt Bias: The Sensitivity of LLM Reasoning to Prompt Variations
Despite the remarkable problem-solving capabilities of Large Language Models (LLMs), their reasoning is highly sensitive to subtle changes in prompt phrasing. As demonstrated in Figure 8, even a single word variation in the prompt can significantly alter the reasoning process and final decision-making. This phenomenon underscores a critical limitation in LLM-based strategic reasoning: models do not inherently generalize optimal strategies but instead rely on heuristic cues embedded within the prompt.
D.1.1 Impact of Word Choice on Strategic Reasoning
D.1.2 Implications for Robust Prompting
This stark contrast highlights the fundamental issue that LLMs do not inherently retrieve the most effective strategic reasoning but are instead disproportionately influenced by linguistic cues. The reliance on explicit terminology for optimal reasoning raises concerns about robustness, as different wordings of the same task can lead to dramatically different problem-solving approaches. This suggests that ensuring reliable strategic reasoning in LLMs requires more than just fine-tuned prompts; it necessitates methods that encourage models to autonomously retrieve and apply domain knowledge without over-reliance on explicit wording cues.
These observations motivate our approach in DReaMAD, where we systematically refine LLMs’ strategic reasoning by structuring prior knowledge retrieval and diversifying input perspectives. By mitigating the sensitivity to prompt variations, our method enhances the robustness and consistency of LLM decision-making across different strategic environments.
Appendix E Ablation study & setup
E.1 Cost effectiveness of DReaMAD
While DReaMAD requires a single model, its inference involves additional prompt steps (e.g., prior knowledge elicitation) and a debate process. However, we believe this test-time scaling method is much more efficient than train-time scaling. As we utilized language models through an API, it was challenging to perform a precise quantitative comparison (e.g., GPU usage time) between our test-time scaling approach and traditional model training. Therefore, we compared the costs using dollar amounts, specifically contrasting the API cost per single game using our method versus the costs incurred when OpenAI fine-tuning models on constructed datasets for NIM-N games.
The NIM-N dataset is constructed by mixing data from three different variants of the NIM game. In all variants, the game starts with 31 stones remaining; however, the rules differ in terms of the maximum number of stones that can be removed per turn—3, 4, or 5, respectively. For each variant, eight distinct game states were sampled and the corresponding optimal action was used as the label, yielding a total of 24 training examples.
The results of this comparison are illustrated in the table above. When fine-tuning a model using API-based fine-tuning (GPT-4o-mini), the performance gradually improved with additional training epochs, achieving win-rates of 0.253, 0.300, 0.420, and 0.460 at 1, 2, 3, and 4 epochs respectively, with corresponding API costs of \$0.013, \$0.023, \$0.033, and \$0.043 (Here, the cost of constructing dataset is not included).
In contrast, our proposed method, DReaMAD, achieved significantly higher performance (0.966) with substantially lower API costs (\$0.0098). These results strongly suggest that our approach not only outperforms traditional fine-tuning methods but also is far more cost-efficient. All the price is calculated by the pricing policy: https://openai.com/api/pricing/
@@ E.1 Cost effectiveness of DReaMAD
E.2 Applicability of Diverse Amplication on Self-Reflection and Self-Consistency
We show whether other self-correction methods—such as Self-Refinement and Self-Consistency—can benefit from structured guidance that enhances reasoning diversity. In DReaMAD, this is operationalized through the Strategic Prior Knowledge Elicitation (SPKE) module, which prompts the model to reinterpret the problem and formulate general strategies before engaging in debate.
To isolate SPKE’s impact, we evaluate DReaMAD, which includes SPKE but excludes debate (see Table 3). We further apply SPKE to Self-Refinement and Self-Consistency and compare them to their vanilla versions. The results show that SPKE alone consistently improves performance across settings.
@@ E.2 Applicability of Diverse Amplication on Self-Reflection and Self-Consistency
E.3 Experiments of Figure 3 setup
For Figure 3, we explain the situation used in this experiment. Figure 9 illustrates a Nim game scenario with a pile of 5 items remaining. On the left (Current State), we present the basic setting and the task: the player (Agent 1) must decide how many items to take given the rules of the Nim variant. Below it (Strong Consistency), we see a single-agent reasoning process where the agent internally evaluates the outcome of different moves and arrives at a conclusion (taking 2 items, leaving 3 to the opponent).
On the right (Multi-Agent Debate), we show a contrasting approach in which two agents (Agent 1 and Agent 2) engage in a debate. Each agent proposes a move and justifies why it would be advantageous. For example, Agent 1 reasons that taking 2 items leaves the opponent with a position that is favorable for Agent 1 (which is wrong reasoning), while Agent 2 counters by proposing to take 1 item for a different strategic benefit (correct reasoning).
Figure 9: The situation and debate process of the experiments in Figure 3

### v5

@@ 2. On your turn, you can remove:
#Task: Based on the current state of the game, decide which pin(s) you will take on this turn.
@@
2. Coordinates use zero-based indices [row, col]. Row 0 is the top row; Col 0 is the leftmost column. Valid ranges: row $\in$ [0, board_height - 1], col $\in$ [0, board_width - 1].
3. From the current position [r, c] the queen may move to **one** of: (a) left: [r, c’] with c’ < c; (b) down: [r’, c] with r’ > r; (c) left-down diagonal: [r + d, c - d] with d > 0.
@@
#Current State: Current position: [row = {r}, col = {c}]. #Task: Based on the current state, decide the next move [row, col].
Table 9: Game state input prompts.
A.2.2 Reasoning and DReaMAD Prompts
Standard, ReAct & CoT Prompts.
In our evaluation of LLMs within the MetaNIM Arena, we compare several key prompting techniques: Standard, Zero-shot Chain-of-Thought (CoT), ReAct Prompting and DReaMAD as shown in Table 8. The distinction between these approaches significantly impacts the model’s reasoning and decision-making process.
Prompting Strategy for Opponent Modeling.
@@ A.2.2 Reasoning and DReaMAD Prompts
Appendix B Ablation study
(a) GPT-4o Wrong Belief Good Belief
(20, 19) (12, 4) (7, 4) (15, 10) (16, 8) (7, 7)
Standard 0.700 0.675 0.600 0.525 0.725 0.850
+ After MAD 0.900 0.750 0.700 0.750 0.750 0.900
(b) GPT-4o-mini Wrong Belief Good Belief
(18, 4) (12, 6) (10, 4) (15, 10) (7, 2) (15, 2)
Standard 0.700 0.875 0.600 0.975 0.950 0.975
+ After MAD 0.850 0.950 0.750 1.000 0.950 1.000
(c) Gemini-1.5-pro Wrong Belief Good Belief
(20, 19) (12, 4) (4, 4) (15, 10) (10, 4) (16, 8)
Standard 0.650 0.600 0.600 0.800 0.625 0.675
+ After MAD 0.700 0.650 0.800 0.800 0.800 0.700
(d) Gemini-1.5-flash Wrong Belief Good Belief
(12, 6) (12, 4) (7, 4) (15, 4) (20, 19) (7, 7)
Standard 0.800 0.700 0.525 0.750 0.500 0.750
+ After MAD 0.850 1.000 0.750 0.750 0.650 1.000
Table 10: Belief entrenchment across models: showing that even after the debate concludes, strongly consistent actions continue to exhibit strong consistency, reinforcing initial static belief in the Fibonacci game. A wrong belief occurs when the model’s biased response deviates from the optimal action, while a correct belief refers to cases where the wrong dominant response aligns with the optimal action. (a, b, c, d) indicate the state where the remaining items are a, and player can take the items maximum to b.
B.1 Instance of Belief Entrenchment
In NIM: In Figure 9, we find that MAD amplifies models’ pre-existing biases rather than refining their reasoning. To investigate this, we first identify game states where each model exhibits strong consistency, i.e., consistently selecting the same action across multiple trials. For each such state, we conduct MAD using two identical agents instantiated from the same model. Each agent generates 20 responses (40 per state) at a fixed temperature of 0.7, maintained throughout the debate. Initial action distributions (light red) are compared against post-debate distributions after three rounds (blue). If MAD functioned as a self-correction mechanism, we would expect the distribution to shift toward the optimal action.
However, we found the opposite: regardless of whether the initial reasoning was correct, the debate process consistently amplifies pre-existing beliefs rather than mitigating them. For instance, in Figure 9 (top-left), GPT-4o-mini initially selects a suboptimal action (Action 3) 82.5% of the time. After the debate, this frequency increases to 90.0%, while the proportion of optimal responses drops further. Rather than correcting errors, the debate reinforces strongly consistent—but incorrect—responses. This pattern persists across model families. The Gemini models (bottom row of Figure 9) exhibit similar behavior, regardless of whether the initial belief aligns with optimal play (“good belief”) or not (“wrong belief”). In both cases, MAD strengthens the dominant trajectory without introducing new strategic insight.
Figure 2 provides a detailed view: two LLM agents receive the same input and begin debating. Despite initial divergence, they converge quickly—after the first round—on a shared line of reasoning. Crucially, this convergence occurs even when the initial consensus is incorrect, illustrating that MAD often serves as an amplifier of belief rather than a correction mechanism.
Figure 9: Belief entrenchment in NIM game by MAD. We compared initial action distribution and action distribution after 3 rounds of debates.
In Fibonacci: Extending our NIM analysis, we evaluate MAD’s effect in the Fibonacci game, a more complex setting with move constraints and dynamic interactions. As in NIM, we identify states exhibiting strong consistency and categorize them into two groups: those where consistent responses align with the optimal strategy, and those where they do not. We then apply MAD to examine how response distributions evolve post-debate. Consistent with the NIM results, MAD reinforces the model’s initial beliefs in Fibonacci as well. As shown in Table 10, the frequency of initially consistent actions increases after debate, regardless of whether those actions are optimal. On average, reinforcement rises by 9.17% in GPT models and 12.29% in Gemini models, indicating a systematic amplification of dominant reasoning patterns across architectures.
@@ B.1 Instance of Belief Entrenchment
B.2 Cost effectiveness of DReaMAD
While DReaMAD requires a single model, its inference involves additional prompt steps (e.g., prior knowledge elicitation) and a debate process. However, this test-time scaling method is much more efficient than train-time scaling. As we utilized language models through an API, we compared the costs using dollar amounts, contrasting the API cost per single game under our method with the costs of OpenAI fine-tuning models on constructed datasets for NIM-N games.
The NIM-N dataset is built from three different variants of the NIM game. Each game starts with 31 stones, but the maximum removable stones per turn differs (3, 4, or 5). For each variant, eight distinct game states were sampled with their optimal actions as labels, yielding 24 training examples in total.
The results of this comparison are illustrated in the table above. When fine-tuning a model using API-based fine-tuning (GPT-4o-mini), the performance gradually improved with additional training epochs, achieving win-rates of 0.253, 0.300, 0.420, and 0.460 at 1, 2, 3, and 4 epochs respectively, with corresponding API costs of \$0.013, \$0.023, \$0.033, and \$0.043 (Here, the cost of constructing dataset is not included). In contrast, our proposed method, DReaMAD, achieved significantly higher performance (0.966) with substantially lower API costs (\$0.0098). These results strongly suggest that our approach not only outperforms traditional fine-tuning methods but also is far more cost-efficient. All prices are calculated by the pricing policy: https://openai.com/api/pricing/
@@ B.2 Cost effectiveness of DReaMAD
Figure 10: Reasoning comparison based on the presence of the keyword ‘Fibonacci’ in the prompt. Explicit usage aligns reasoning with optimal strategy, demonstrating the impact of lexical cuing on strategic planning.
B.3 Applicability of Diverse Amplification on Self-Reflection and Self-Consistency
We show whether other self-correction methods, such as Self-Refinement and Self-Consistency, can benefit from structured guidance that enhances reasoning diversity. In DReaMAD, this is operationalized through the Strategic Prior Knowledge Elicitation (SPKE) module, which prompts the model to reinterpret the problem and formulate general strategies before engaging in debate. To isolate SPKE’s impact, we evaluate DReaMAD, which includes SPKE but excludes debate (see Table 1). We further apply SPKE to Self-Refinement and Self-Consistency and compare them to their vanilla versions. The results show that SPKE alone consistently improves performance across settings.
B.4 Belief Entrenchment in NIM Scenario
For Figure 4, we describe the specific scenario where debate facilitates belief entrenchment. The setup is a NIM-Normal game with 5 items remaining. An agent must decide how many items to take. With Strong Consistency, an agent evaluates possible moves on its own and concludes that taking 2 items (leaving 3 for the opponent) would lead to a winning position. Conducting MAD, two agents exchange their reasoning before deciding. An agent suggests taking 2 items but provides flawed reasoning, while another argues that taking 1 item is better since it leaves the opponent with 4 items, a multiple of 4, which is known to be strategically advantageous.
B.5 A Word Change in Prompts Leads to Different Output